OpenStudy (ray1998):

What are the real or imaginary solutions of the polynomial equation? x^4 - 52x^2 + 576 = 0 A. 4, -4 B. 4, -6 C. 4, -4, 6, -6 D. No solution Fan and medal will be given out to whoever is most helpful. Thank you in advance!!! :D

OpenStudy (er.mohd.amir):

compare with ax^2+bx+c=0

OpenStudy (michele_laino):

this is a bi-quadratic equation, and in order to solve it, it is necessary to make this variable change: \(x^2=z\), so we get: \(z^2-52z+576=0\) so, please solve such equation for \(z\)

OpenStudy (er.mohd.amir):

or use 52=36+16 and 36*16=576

OpenStudy (ray1998):

Thank you guys! :) this was helpful.

OpenStudy (anonymous):

Are u with me

OpenStudy (anonymous):

Let's go

OpenStudy (anonymous):

If Y=x² substitute Y into X² Y²-52Y+576=0

OpenStudy (anonymous):

Now picking from Mohd.Amir clue Y²–16Y–36Y+576=0

OpenStudy (anonymous):

Do u know one thing I hate about maths is the big number but thanks to technology let's move let's move to the next level by finding Y even though we are still seeing it

OpenStudy (ray1998):

Oh! Sorry, I'm really busy and got sidetracked. Thank you for all your help :) @Daniel56k

OpenStudy (anonymous):

Y(Y–16)–36(Y–16)=0

OpenStudy (anonymous):

(Y–16)(Y–36)

OpenStudy (ray1998):

The answer is C.

OpenStudy (anonymous):

Okay let me just finish the calculation up please

OpenStudy (anonymous):

Just kidding u are right

OpenStudy (ray1998):

Lol ^.^ thank you

OpenStudy (anonymous):

Y=16,Y=36 Substitution √x²=√16 X=4 X=6

OpenStudy (ray1998):

Yep

OpenStudy (anonymous):

I found the number right

OpenStudy (anonymous):

Yeppeeeeeee

OpenStudy (ray1998):

Lol ^.^

OpenStudy (anonymous):

My head is so hot right now it can burn down a whole house.......bye

OpenStudy (anonymous):

Have to cool my head off with so other maths

OpenStudy (anonymous):

Some other maths